From the top of a lighthouse 60m high, a boat with an angle of depression of 37° If the boat travels linearly 35m towards the tower. What is the new angle of
elevation?

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

From the top of a lighthouse 60m high, a boat with an angle of depression of 37° If the boat travels linearly 35m towards the tower. What is the new angle of
elevation?
-------------------
tan(37) = 60/d where d is the distance from the base of the lighthouse to the boat.
d = 60/tan(37) = ~ 79.62 meters
-----
After moving 35 meters,
tan(A) = 60/(79.62-35)
A = ~ 53.36 degs, the angle of depression from the lighthouse, not angle of elevation.

Answer by Theo(13342)

the new angle of elevtion is 53.36139457 degrees.

see my worksheet below:



let me know if you have any question.

theo